OpenStudy (anonymous):

can anyone help with geo 1 really need help with transformations will give medal

OpenStudy (anonymous):

What do you need?

OpenStudy (anonymous):

A trriangle with vertices at A(0,0) B(0,4) and C(6,0) is dialated to yield a triangle with vertices at A(0,0), B(0,10, and C(15,0). The orgin is the center of dilation. What is the scale factor of the dilation? I think it is 3 but not sure

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

A lot of times it helps to graph the shape first, then have a look at it. Have you done this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

several times

OpenStudy (anonymous):

im thinking i dont understand the concept

OpenStudy (anonymous):

Well a dilation implies that the two triangles will be similar, correct?

OpenStudy (anonymous):

yes just either larger or smaller

OpenStudy (anonymous):

So you can set up an inequality using two of the sides I believe.

OpenStudy (anonymous):

\[\frac{ 6 }{ 15 }=\frac{ 4 }{ 10 }\]

OpenStudy (anonymous):

Where 6 and 4 are your old side lengths, and 10 and 15 are the new lengths after dilation.

OpenStudy (anonymous):

And by reducing both, you can see that you get \[\frac{ 2 }{ 5 }\] for both.

OpenStudy (anonymous):

2/5 =2/5

OpenStudy (anonymous):

so would that be 2.5

OpenStudy (anonymous):

Right, and so the dilation is of the ratio 2:5, or I believe it'd be more properly written as 1:2.5

OpenStudy (anonymous):

my choices were 4.5, 2, 2.5, 3

OpenStudy (anonymous):

Perfect, so 2.5 it is!

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

wow you made it so much easier i have been strugling with this, so to get it right I take the two sides with the same origin and do an inequality. thank u so much

OpenStudy (anonymous):

i wish my teacher told me that

OpenStudy (anonymous):

Haha well it depends on the shape too. This works because those lines were straight so you could find their lengths without doing any real math. If they had been diagonal lines you'd have to use the distance formula. But with similar shapes, you can always set up inequalities like that, it's a neat trick to keep in mind. Good luck!

OpenStudy (anonymous):

Can you help me with another one?